OpenStudy (anonymous):

When you weigh yourself on good old terra firma (solid ground), your weight is 130 lb. In an elevator your apparent weight is 100 lb. What are the direction and magnitude of the elevator's acceleration? I know that the direction is downward because your weight it more than the apparent weight. I know that Wa(apparent weight)=m(g-a) but I am confused on how to incorporate your weight as well?

OpenStudy (anonymous):

I'm confused why they would do this question in lbs and not in kgs. But either way, the 'm' in the formula is mass. Weight = mass * gravity.

OpenStudy (anonymous):

yeah i think that is why i am getting confused. But do I just use the apparent weight or do I have to include calculations of the actual weight?

OpenStudy (amistre64):

130 = mg 100 = ma 130/g = 100/a a = g 100/130 lb/lb cancels and we are left to play with the acceleration in whatever units g is in .

OpenStudy (amistre64):

or am i underthinking this?

OpenStudy (anonymous):

so the masses equal each other?

OpenStudy (amistre64):

of course, the person hasnt changed ... their mass remains relatively constant inside and outside an elevator

OpenStudy (anonymous):

I got 7.546 but it isnt correct.

OpenStudy (amistre64):

i get 24.6 ish, ft/ sec^2

OpenStudy (anonymous):

the acceleration is in m/s^2

OpenStudy (amistre64):

you sure it is? if so then convert 130lbs = 578.26881 N 100 lbs = 444.82216 N

OpenStudy (amistre64):

assuming g = 9.8; a = 7.538 some use g=10 .. what do you spose we should use?

OpenStudy (amistre64):

if we are given lbs i think the answer is in ft/sec^2

OpenStudy (anonymous):

Nope definitely wants it in m/s^2. This is the last question that I have to do and I cannot get it. I honestly have no idea...

OpenStudy (amistre64):

can you attach a picture of it? or a link?

OpenStudy (anonymous):

No picture :/ sorry. Its just a word problem and thats all it gives

OpenStudy (amistre64):

well, converting to Newtons should have done it then A = mg ; m = A/g B = ma ; m = B/a a = g B/A it might seem simplistic, but a diagram might help? dunno

OpenStudy (amistre64):

did you include the negative? or is there a different downward indicator?

OpenStudy (amistre64):

and for the record, i get 7.538 using lbs and newtons. either way 9.8 (100/130) = 9.8 (444.82216 / 578.26881) = 7.538...

OpenStudy (anonymous):

but it is going down and isnt acceleraton positive when going downward.

OpenStudy (anonymous):

Idk the program says that it answer differs by more than 100% so maybe it is negative?

OpenStudy (amistre64):

acceleration is negative (by convention/tradition) when making an equation for motion, we have to start with a negative acceleration to suggest that its pulling us downward.

OpenStudy (amistre64):

spose we toss a ball upwards a = -g v = -gt + k d = -1/2 g t^2 +kt +c for some given initial upward velocity, k, and initial height, c.

OpenStudy (anonymous):

but the elevator is moving downwards so a=g right?

OpenStudy (amistre64):

its accelerating downwards, so the acceleration is negative: its a vector pointing down.

OpenStudy (amistre64):

if a=g, then we free fall

OpenStudy (anonymous):

oh okay. So a=-7.538

OpenStudy (amistre64):

if we let g=9.8, then yes.

OpenStudy (amistre64):

see if there is some "decimal" rnding place they want. rnd to 2 decimals or whatever

OpenStudy (anonymous):

yeah it was wrong and it was my last submission but thank you for your help

OpenStudy (amistre64):

lets consider some stuff the elevator is eating up some acceleration, so it may be that the elevator is g-a 9.8 - 7.538 = 2.262 does this make sense? if the elevator is doing 9.8, then the acceleration we feel is 0, free fall

OpenStudy (amistre64):

9.8 - (0) = 9.8 (elevator) if the elevator is not moving 9.8 - (9.8) = 0 (elevator) soo, 9.8, minus our apparent acceleration, is equal to the elevators acceleration.

OpenStudy (irishboy123):